Please this question is hard help me

Three point charges with values q_1 = 8.00 nC, q_2 = 3.00 nC, and

q_3 = 7.00 nC are placed on three consecutive corners of a square whose side measuress = 4.00 m. Point A lies at the fourth corner of the square, diagonally across from q_2. Point B lies at the center of the square.A.) What is the electric potential at point A?
Answer #1
Three point charges with values q1 = 6.00 µC, q2 = 3.00 µC, and q3 = 5.00 µC are placed on three consecutive corners of a square whose side measures a =3.00 m. Point A lies at the fourth corner of the square, diagonally across from . Point B lies at the center of the square.

The length of the diagonal = 3 * v2

A)What are the values of the electric potential at point A?

The electric potential created by a point charge q, at a distance r from the charge (relative to the potential at infinity), can be shown to be
Ve = k * q ÷ r)
k = Coulomb’s constant = 9 * 10^9 N * m^2/Coulomb^2
q = charge in Coulombs
r = distance in meters
The electric potential due to a system of point charges is equal to the sum of the point charges' individual potentials. This fact simplifies calculationssignificantly, since addition of potential (scalar) fields is much easier than addition of the electric (vector) fields.

Distance from q1 to A =
Distance from q2 to A = length of diagonal
Distance from q3 to A =

V1 = (9 * 10^9) * (8 * 10^-6) ÷ d1=a
V2 = (9 * 10^9) * (3 * 10^-6) ÷ d2 =b
V3 = (9 * 10^9) * (7 * 10^-6) ÷ d3 =c

Total Electric potential at A=a+b+c
answered by: LOP